OpenStudy (anonymous):

I will give a medal who answers this and fan them It takes Brian 15 hours longer to build a model car than it takes John. If they work together, they can build the model car in 4 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Brian to build the car on his own

jimthompson5910 (jim_thompson5910):

x = time it takes for Brian to build the model car on his own y = time it takes for John to build the model car on his own we're told that `It takes Brian 15 hours longer to build a model car than it takes John` so x = y+15 \[\large \frac{1}{\text{Brian's Time}}+\frac{1}{\text{John's Time}} = \frac{1}{\text{Time if they work together}}\] \[\large \frac{1}{x}+\frac{1}{y} = \frac{1}{4}\] \[\large \frac{1}{y+15}+\frac{1}{y} = \frac{1}{4}\] Do you know how to solve for y from here?

OpenStudy (anonymous):

No im so sorry this isn't my strongest subject.

jimthompson5910 (jim_thompson5910):

what is the LCD of all the fractions?

OpenStudy (anonymous):

Wouldn't it be 1?

jimthompson5910 (jim_thompson5910):